OpenStudy (anonymous):

f(x)=-2x^2+1 and g(x)=3, evaluate f(x) devided by g(x) for x=2

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (misssunshinexxoxo):

First, what do you think the answer is?

OpenStudy (anonymous):

not exactly sure

OpenStudy (anonymous):

@abb0t ;D

OpenStudy (anonymous):

pro-3/2 ?

OpenStudy (anonymous):

-3/2 *

OpenStudy (anonymous):

?

Mehek (mehek14):

first put 2 for x \(2^2=4\) \(-2*2 + 1 = -3\) so \(f(x)=-3\) now divide that by g(x) \(\dfrac{-3}{3}=-1\)

Mehek (mehek14):

@DUDE..IM..A..DUCK